need NSWER FOR THE SECOND ONE THANKYOU
Answer #1

2.

We have following information about Ensong Corp.

WACC(k) = 9.5%

FCFF grow for next two years(G1) = 20%

FCFF grow indefinitely after 2 year(G2) = 4%

Free cash flow to Firm = $ 181 millions .. (for calculation please refer to below spreadsheet)

In order to calculate Price Per Share of Ensong Corp., We first calculate the Value of Firm using FCFF.

FCFFi FCFF2FCFFn TerminalV alue V alueof Firm = n+1

Where,

k = Weighted average cost of capital

n= year

Terminal Value = PV of constant growing FCFF

FCFF TerminalValue k -g

where,

FCFF = Free cash flow to firm growing at constant rate of growth

g = growth rate

k = wacc

Calculation of Value of Firm

Please refer below sheet.

Calculation of Value of Firm WACC G1 G2 9.50% 20% 4% in $ milions PVF(9.5% 0.91324201 0.83401097 0.76165385 Year FCFF PV 217.20 260.64 271.07 198.36 2 217.38 3 Terminal Value 4928.47 0.76165385 3753.78 Value of Firm 4169.52

Formula reference -

Calculation of Value of Firm WACC G1 G2 0.095 0.2 0.04 (in $milions PVF(9.5% -1/(1+$Q$3)AP7 -1/(1+$Q$3)AP8 -1/(1+SQ$3)P9 1/(1+$Q$3)AP10 Year FCFF PV -181 (1+04 -Q7 (1+04 -Q8*1.04 Terminal Value 09/(03-05) R10 S10 SUM(T7:T10 Value of Firm

Thus, Value of Firm = $ 4,169.52 millions.

Now, Assuming the value of Debt and Preferred stock in balance sheet is market value of them

Therefore, Value of common Equity = Value of Firm - Debt - Preferred stock

= 4,169.52-963-10

= $ 3,196.52 millions

Price Per Share = Value of common equity/no. of common stock

= 3,196.52/100

= $ 31.97

for calculation of free cash flow please refer to below spreadsheet.

M13 C D Information from Balance Sheet of Ensong Co In $ millions Total Fixed Assets Current Assets Current Liabilities Long Term Debt retained 100% of its Income as Income statement not provided 2015 2016 1530 1802 938 1011 544 591 937 963 987 1259 Assuming Ensong co 2016 In $ millions 272 181 453 204 657 Net Income (EAT Add Tax | EBT Add Interest (EBIT-EBT -|EBIT EBIT Depreciation Retained Earnin Tax Rate Cash Non-cashWorking Capital 376 399 657 197261 777 1049 40% 21 10 Computation of Free Cash Flow of Esnsong Co 2016 12 13 14 15 16 17 18 19 20 21 657 181 64 23 336 181 EBIT 18 Less Tax Add Depreciation Less Increase in non-cash working Capital Less Capital Expenditure Free Cash Flow (FCF Computation of Return on Invested Capital (ROIC 2016 Invested Capital (opening Balance) Net Operationg Profit after tax (NOPAT) ROIC 1924 394 20.49% 23 24